# K-theory and localizing invariants of large categories

### v2

@@ 5.4. Localizing invariants of categories of cosheaves on continuous posets
Let now be a copresheaf of dualizable categories on for By a -valued copresheaf on we mean a section of the associated biCartesian fibration over We denote by this category of copresheaves. Further, we denote by the full subcategory of cosheaves: it consists of copresheaves such that We denote by the cosheafification functor, i.e. the right adjoint to the inclusion.
Again, we denote by the left Kan extension functor, given by
@@
1) The category is dualizable.
2) For each compact element the functor takes values in cosheaves, hence we have a fully faithful strongly continuous functor
3) The functors induce an isomorphism
@@
It follows that the category is dualizable. This proves 1).
Part 2) is clear from the definitions.
We prove 3). Consider the short exact sequence
@@
It follows by adjunctions that the functors are fully faithful () and their images are semi-orthogonal as required. Since the functors take values in cosheaves, we obtain the semi-orthogonal decomposition (5.8).
@@ 6.1. Sheaves on finite CW complexes
The following result describes arbitrary (accessible) localizing invariants of categories of sheaves on finite CW complexes, in the case of constant coefficients. Here we consider finite CW complexes both as topological spaces and as -groupoids (i.e. objects of -category of finite spaces).
@@
Therefore we have a well defined functor such that for a finite CW complex we have Since it suffices to prove that commutes with pullbacks.
This is straightforward. Namely, consider cellular closed embeddings of finite CW complexes Then both pullback functors and are strongly continuous localizations and we have a pullback square
@@ 6.2. Sheaves and K-sheaves on locally compact Hausdorff spaces
Let now be a locally compact Hausdorff space, and consider a presheaf on with values in In particular, we require that the restriction functors are strongly continuous.
We will consider -valued (pre)sheaves on Recall that is a sheaf if and only if it satisfies the following conditions:
for any open we have a pullback square
for any open the map is an isomorphism.
@@
Suppose that is a presheaf which satisfies descent for finite covers, i.e. the conditions (i) and (ii). Then is a sheaf. Equivalently, the natural map is an isomorphism.
In particular, the inclusion functor is continuous.
@@
Assuming that satisfies finite descent, it suffices to show that satisfies finite descent.
Clearly, so (i) holds. To see that (ii) holds for we first introduce the following notation: for an open we denote by the poset Now take some open subsets and consider the maps
The standard arguments show that both maps (6.1) and (6.2) are cofinal (hence, the maps between the same posets with the opposite order are final). This directly implies that (ii) holds for Thus, is indeed a sheaf.
For the final assertion of the proposition, it suffices to notice that the class of presheaves satisfying finite descent is closed under colimits. ∎
We consider the category as the first approximation of the category Note that if is a presheaf with values in then we know the finitary localizing invariants of the category by Theorem 5.4. To construct the further approximations, it is convenient to use the language of -sheaves, which we know recall.
We first extend the presheaf to locally closed subsets by left Kan extension: for a locally closed we put
where runs over open subsets such that is a closed subset of In particular, we obtain a presheaf of presentable stable categories on
Denote by the poset of compact subsets of
@@
1) Denote by the category of -presheaves, i.e. the category of -valued presheaves on
2) We denote by the full subcategory of continuous -presheaves, i.e. -presheaves such that for any compact we have an isomorphism
Here if contains an open neighborhood of
3) We denote by the full subcategory formed by continuous -presheaves satisfying additional conditions:
i)
ii) For compact subsets the following is a pullback square:
Note that the poset is continuous, and iff Hence, the category is the category of -valued cosheaves for the topology from Subsection 5.1.
The following is an (almost) straightforward generalization of Lurie’s theorem for the case of a constant presheaf
@@
[Lur09, Theorem 7.3.4.9, Corollary 7.3.4.10] 1) Consider the following functors
They are quasi-inverse equivalences.
@@
Part 1) is proved by the same argument as [Lur09, Theorem 7.3.4.9], and in fact we do not need the strong continuity of the restriction functors
We prove 2). Let be a -valued sheaf, and denote by the associated continuous -presheaf. Clearly, Consider compact subsets and put It suffices to show that the square
@@
over open neighborhoods
Clearly, we have
@@
(here we used that the restriction functors are strongly continuous). The proof for the remaining two objects of the square (6.5) is similar.
The proof in the other direction is as in [Lur09, Theorem 7.3.4.9] (and here we do not need the strong continuity of the restriction functors). ∎
@@
Let be a directed poset, and consider a functor Consider a section of the associated biCartesian fibration over Denote by the transition functors, and denote by the functors to the colimit. Then for any we have
@@
We consider as a poset with the reverse inclusion order.
Consider the set of all collections of compact (=closed) subsets of of the form as in (6.7). Define the partial order on by setting if Then is a directed poset.
@@
1) If is locally compact, then the poset has a natural Grothendieck pretopology, where the family is a cover if is finite and Then for each compact the poset is a cofinal family of covering sieves for
2) If is compact and is a sieve as above, then we have a final functor
@@
Consider the full subcategory of presheaves satisfying only the finite descent, i.e. sheaves for the finitary topology on described in Remark 6.6 1). Then the corresponding finitary sheafification functor is given by (6.8). Thus, we need to show the following: for the -presheaf is continuous. This is straightforward. ∎
@@
1) For each the functor is a strongly continuous localization. The essential image of the right adjoint consists of continuous presheaves such that for each compact we have
@@
Observe that the analogue of part 1) for categories holds trivially. To prove the statement 1), we only need to show the following.
Claim. Let be a -presheaf satisfying (6.9). If each of the pullbacks is in then also
@@
We may assume that i.e. Take a compact subset It suffices to show that the map
is an isomorphism, and similarly for and
@@
We need to show that this condition holds if and only if is a -sheaf.
The “if” direction is obvious. We show the “only if” direction. Take some pair of compact subsets For any open neighborhoods, denote by the sieve Applying the condition (6.10) to and we see that the following is a pullback square:
@@ 6.3. Localizing invariants of categories of sheaves
Suppose that is compact. For a sieve the natural functor
@@
This is basically an iterated application of Propositions 4.11 and 6.9. We prove the statement by induction on where the case is clear.
Let where We put
@@
are equivalences.
@@
Let be locally compact Hausdorff, and a presentable stable category. Given a -presheaf we define another -presheaf by the formula
where runs over open-closed subsets of We have a natural map
1) If is continuous, then so is
2) Assuming that is continuous, the map of -sheaves is an isomorphism.
@@
is a bijection of sets.
2) Recall that is the sheafification of for the finitary topology on Consider the weaker pretopology on for which a family is a covering if is finite and Denote by the corresponding sheafification functor. It suffices to show that the map is an isomorphism.
Let us observe that the weak sheafification functor is given by
@@
and the fact that the functor is strongly continuous.
Now observe that we have a generating collection of sheaves
@@
and by the above discussion each map is compact. This shows that the category is dualizable.
To compute we first reduce to the case when is compact. Denote by the one-point compactification of and denote by the embedding. Let be the presheaf of dualizable categories on obtained by extension by zero of Then hence we may replace the pair with the pair Thus, we may and will assume that is compact.
@@ 6.4. Compact objects
The implication is clear.
. Suppose that is compact.
We first show the following: if for some then there is an open neighborhood such that Indeed, since and is compact, we deduce that is a direct summand of some sheaf of the form where Hence the restriction of to vanishes.
Now let be a point such that Since the functor is strongly continuous, the object is compact in Hence, there is an open neighborhood and a compact object such that Shrinking if necessary, we can lift the identity map of to a map in the category Denote by the associated map of sheaves. Choosing an open such that we see that is a compact object of whose stalk at vanishes. By the above, we deduce that vanishes in a neighborhood of Hence, shrinking further, we may assume that as required.
Finally, to see that has a compact support, note that Since is compact, it is a direct summand of some sheaf which has compact support. ∎
@@
The category is compactly generated, and is locally profinite.
@@
Then each sheaf in is compact and the right orthogonal to consists of sheaves such that for every open-compact Since is locally profinite, this means that Hence, generates the category
Choose a point The pullback to is a strongly continuous quotient functor Since the category is compactly generated, so is
Next, we reduce to the case when is compact. If not, consider the short exact sequence
Since the categories on the left and on the right are compactly generated, so is the category in the middle. Hence, we may and will assume that is compact.
Suppose that is not profinite. Choose a connected component such that is not a single point. Then is closed, hence compact Hausdorff. The functor is a strongly continuous localization. Since the category is compactly generated, so is the category Replacing with we may and will assume that is connected.
Choose a point Choose a non-zero compact object and consider the skyscraper sheaf on Then we can find a directed system of compact -valued sheaves such hat By Proposition 6.14 each is a locally constant sheaf with compact stalks.
Since is compact in we can find and a splitting Using compactness of again, we can find an open neighborhood and a map which lifts Choose a compact such that is a neighborhood of Let be the connected component of in Then hence we may and will assume that
So we have the map Take some Since the sheaf is locally constant, the set
@@ Appendix A Monomorphisms and epimorphisms of presentable and dualizable categories
Here we show that the monomorphisms and epimorphism in and are as expected: these are respectively the fully faithful functors and the quotient functors.
@@
Let be a continuous functor between presentable stable categories.
1) is a monomorphism in if and only if is fully faithful.
2) is an epimorphism in if and only if is a quotient functor.
@@
is a monomorphism, hence is fully faithful (since and are stable).
2) Again, the “if” part is clear. For the “only if” part, note that where both functors are Hence, the functor is an equivalence, in particular, it is fully faithful. Hence, for any the square
@@
Let be a strongly continuous functor between presentable stable categories.
1) is a monomorphism in if and only if is fully faithful.
2) is an epimorphism in if and only if is a quotient functor.
@@
1) The “if” part is clear. We prove the “only if” part.
Suppose that is a monomorphism. Then for a presentable stable category and strongly continuous functors we have By adjunction, we have
@@
hence which exactly means that is fully faithful. ‘
2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎
@@
Let be a strongly continuous between dualizable categories.
1) is a monomorphism in if and only if is fully faithful.
2) is an epimorphism in if and only if is a quotient functor.
@@
1) The “if” direction is clear. We prove the “only if” direction.
Suppose that a functor is a monomorphism in By Proposition 1.88 the functor is fully faithful. Hence, the functor is fully faithful. Arguing as in the proof of Proposition A.1, we conclude that is fully faithful.
2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎

### v3

@@ 5.4. Localizing invariants of categories of cosheaves on continuous posets
Let now be a copresheaf of dualizable categories on for By a -valued copresheaf on we mean a section of the associated cocartesian fibration over (which is also cartesian). We denote by this category of copresheaves. Further, we denote by the full subcategory of cosheaves: it consists of copresheaves such that We denote by the cosheafification functor, i.e. the right adjoint to the inclusion. Again, we denote by the left Kan extension functor, given by
@@
1) The category is dualizable. 2) For each compact element the functor takes values in cosheaves, hence we have a fully faithful strongly continuous functor 3) The functors induce an isomorphism
@@
It follows that the category is dualizable. This proves 1). Part 2) is clear from the definitions. We prove 3). Consider the short exact sequence
@@
It follows by adjunctions that the functors are fully faithful () and their images are semi-orthogonal as required. Since the functors take values in cosheaves, we obtain the semi-orthogonal decomposition (5.8). As in the proof of Theorem 5.4, we see that after applying to the sequence (5.7) we get a split exact triangle:
This proves 3). ∎
@@ 6.1. Sheaves on finite CW complexes
The following result describes arbitrary (accessible) localizing invariants of categories of sheaves on finite CW complexes, in the case of constant coefficients. Here we consider finite CW complexes both as topological spaces and as -groupoids (i.e. as objects of the -category of finite spaces).
@@
Therefore we have a well defined functor such that for a finite CW complex we have Since and it suffices to prove that commutes with pullbacks. This is straightforward. Namely, consider cellular closed embeddings of finite CW complexes Then both pullback functors and are strongly continuous localizations and we have a pullback square
@@ 6.2. Sheaves and K-sheaves on locally compact Hausdorff spaces
Let now be a locally compact Hausdorff space, and consider a presheaf on with values in In particular, we require that the restriction functors are strongly continuous. We will consider -valued (pre)sheaves on Recall that is a sheaf if and only if it satisfies the following conditions:
for any open we have a pullback square in
for any open the map is an isomorphism in
@@
We call a presheaf reduced if i.e. if the condition (i) holds for Denote by the full subcategory of reduced presheaves. The inclusion functor has a left adjoint where
Note that if is continuous, then so is
Remark 6.2.
In what follows, we crucially use the following property of the continuous lattice if and then In other words, the way below relation in is multiplicative in the terminology of [GHKLMS80].
@@
Assuming that satisfies finite descent, it suffices to show that satisfies finite descent. Clearly, so (i) holds. To see that (ii) holds for we first introduce the following notation: for an open we denote by the poset Now take some open subsets and consider the maps
The standard arguments show that both maps (6.2) and (6.3) are cofinal (hence, the maps between the same posets with the opposite order are final). This directly implies that (ii) holds for Thus, is indeed a sheaf. For the final assertion of the proposition, it suffices to notice that the class of presheaves satisfying finite descent is closed under colimits. ∎
We consider the category as the first approximation of the category Note that if is a presheaf with values in then we know the finitary localizing invariants of the category by Theorem 5.4. To construct the further approximations, it is convenient to use the language of -sheaves, which we now recall. We first extend the presheaf to locally closed subsets by left Kan extension: for a locally closed we put
where runs over open subsets such that is a closed subset of In particular, we obtain a presheaf of presentable stable categories on Denote by the poset of compact subsets of
@@
1) We denote by the category of -presheaves, i.e. the category of -valued presheaves on 2) We denote by the full subcategory of continuous -presheaves, i.e. -presheaves such that for any compact we have an isomorphism
Here if contains an open neighborhood of 3) We denote by the full subcategory formed by continuous -presheaves satisfying additional conditions: i) ii) For compact subsets the following is a pullback square:
Note that the poset is continuous, and iff Hence, the category is the category of -valued cosheaves on for the topology from Subsection 5.1. Again, it is convenient to call a -presheaf reduced if The functor is defined as in (6.1). If is a continuous -presheaf, then so is The following is an (almost) straightforward generalization of Lurie’s theorem for the case of a constant presheaf
@@
Part 1) is proved by the same argument as in [Lur09a, Theorem 7.3.4.9], and in fact we do not need the strong continuity of the restriction functors We prove 2). Let be a -valued sheaf, and denote by the associated continuous -presheaf. We need to prove that is a -sheaf. Clearly, Consider compact subsets and put It suffices to show that the square
@@
over open neighborhoods Clearly, we have
@@
(here we used that the restriction functors are strongly continuous). The proof for the remaining two objects of the square (6.6) is similar. Therefore, is a -sheaf. The proof in the other direction is as in [Lur09a, Theorem 7.3.4.9] (and here we do not need the strong continuity of the restriction functors). ∎
@@
Let be a directed poset, and consider a functor Consider a section of the associated cocartesian fibration over Denote by the transition functors, and denote by the functors to the colimit. Then for any we have
@@
We consider as a poset with the reverse inclusion order. Consider the set of all collections of compact (=closed) subsets of of the form as in (6.8). Define the partial order on by setting if Then is a directed poset.
@@
1) If is locally compact, then the poset has a natural Grothendieck pretopology, where the family is a cover if is finite and Then for each compact the poset is a cofinal family of covering sieves for 2) If is compact and is a sieve as above, then we have a final functor Here for a set we denote by the poset of non-empty subsets of
@@
Consider the full subcategory of -presheaves satisfying only the finite descent, i.e. sheaves for the finitary topology on described in Remark 6.7 1). Then the corresponding finitary sheafification functor is given by (6.9). Thus, we need to show the following: for the -presheaf is continuous. This is straightforward. ∎
@@
Suppose that the presheaf is reduced, i.e. 1) For each the functor is a strongly continuous localization. The essential image of the right adjoint consists of continuous presheaves such that for each compact we have
@@
Observe that the analogue of part 1) for categories holds trivially. To prove the statement 1), we only need to show the following. Claim. Suppose that Let be a -presheaf satisfying (6.10). If each of the pullbacks is in then also
@@
We need to show that this condition holds if and only if is a -sheaf. The “if” direction is obvious. We show the “only if” direction. Take some pair of compact subsets For any open neighborhoods, denote by the sieve Applying the condition (6.11) to and we see that the following is a pullback square:
@@ 6.3. Localizing invariants of categories of sheaves
Suppose that is compact and For a sieve the natural functor
@@
This is basically an iterated application of Propositions 4.11 and 6.10. We prove the statement for a sieve by induction on where the case is clear. Let We put
@@
are equivalences. Applying Propositions 6.9 and 6.10, we see that the composition
is a localization (both functors are localizations). It follows from Proposition 4.11 that the pullback square
in is also a pullback square in and applied to this square is a pullback square in This proves the proposition. ∎
For a topological space we denote by the set of open-closed subsets of
@@
Let be locally compact Hausdorff space, and a presentable stable category. Given a reduced -presheaf we define another -presheaf by the formula
We have a natural map 1) If is continuous, then so is 2) Assuming that is continuous, the map of -sheaves is an isomorphism.
@@
is a bijection of sets. 2) Recall that is the sheafification of for the finitary topology on Consider the weaker pretopology on for which a family is a covering if is finite and Denote by the corresponding sheafification functor. It suffices to show that the map is an isomorphism. Let us observe that the weak sheafification functor on reduced -presheaves is given by
@@
and the fact that the functor is continuous. Now observe that we have a generating collection of sheaves
@@
and by the above discussion each map is compact. This shows that the category is dualizable. To compute we first reduce to the case when is compact. Denote by the one-point compactification of and denote by the embedding. Let be the presheaf of dualizable categories on obtained by extension by zero of Then hence we may replace the pair with the pair Thus, we may and will assume that is compact. Also, we may and will assume that Consider the -presheaf given by Then is a continuous -presheaf. By Proposition 6.12, for a compact we have a natural isomorphism
Applying Propositions 6.10, 6.13 and 6.14, we obtain the following chain of isomorphisms
This proves the theorem. ∎
@@ 6.4. Compact objects
The implication is clear. . Suppose that is compact. We first show the following: if for some then there is an open neighborhood such that Indeed, since and is compact, we deduce that is a direct summand of some sheaf of the form where Hence the restriction of to vanishes. Now let be a point such that Since the functor is strongly continuous, the object is compact in By Proposition 1.71 there is an open neighborhood and a compact object such that Shrinking if necessary, we can lift the identity map of to a map in the category Denote by the associated map of -valued sheaves on Choosing an open such that we see that is a compact object of whose stalk at vanishes. By the above, we deduce that vanishes in a neighborhood of Hence, shrinking further, we may assume that as required. Finally, to see that has compact support, note that Since is compact, it is a direct summand of some sheaf which has compact support. ∎
@@
The category is compactly generated and is locally profinite.
@@
Then each sheaf in is compact and the right orthogonal to consists of sheaves such that for every open-compact Since is locally profinite, this means that Hence, generates the category Choose a point The pullback to is a strongly continuous quotient functor Since the category is compactly generated, so is Next, we reduce to the case when is compact. If not, consider the short exact sequence
in Since the categories on the left and on the right are compactly generated, so is the category in the middle (by Proposition 3.3). Hence, we may and will assume that is compact. Suppose that is not profinite. Choose a connected component such that is not a single point. Then is closed, hence compact Hausdorff. The functor is a strongly continuous localization. Since the category is compactly generated, so is the category Replacing with we may and will assume that is connected. Choose a point Choose a non-zero compact object and consider the skyscraper sheaf on Then we can find a directed system of compact -valued sheaves such hat By Proposition 6.15 each is a locally constant sheaf with compact stalks. Since is compact in we can find and a splitting Using compactness of again, we can find an open neighborhood and a map which lifts Choose a compact such that is a neighborhood of Let be the connected component of in Then hence we may and will assume that So we have the map Take some Since the sheaf is locally constant, the set
@@ Appendix A Monomorphisms and epimorphisms of presentable and dualizable categories
Here we show that the monomorphisms and epimorphisms in and are as expected: these are respectively the fully faithful functors and the quotient functors.
@@
Let be a continuous functor between presentable stable categories. 1) is a monomorphism in if and only if is fully faithful. 2) is an epimorphism in if and only if is a quotient functor.
@@
is a monomorphism, hence is fully faithful (since and are stable). 2) Again, the “if” part is clear. For the “only if” part, note that where both functors are Hence, the functor is an equivalence, in particular, it is fully faithful. Hence, for any the square
@@
Let be a strongly continuous functor between presentable stable categories. 1) is a monomorphism in if and only if is fully faithful. 2) is an epimorphism in if and only if is a quotient functor.
@@
1) The “if” part is clear. We prove the “only if” part. Suppose that is a monomorphism. Then for a presentable stable category and strongly continuous functors we have By adjunction, we have
@@
hence which exactly means that is fully faithful. ‘ 2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎
@@
Let be a strongly continuous between dualizable categories. 1) is a monomorphism in if and only if is fully faithful. 2) is an epimorphism in if and only if is a quotient functor.
@@
1) The “if” direction is clear. We prove the “only if” direction. Suppose that a functor is a monomorphism in By Proposition 1.87 the functor is fully faithful. Hence, the functor is fully faithful. Arguing as in the proof of Proposition A.1, we conclude that is fully faithful. 2) follows from Proposition A.1 since the functor is conservative and commutes with colimits. ∎